Calculate the mass of chlorine gas contained in a 5.00 dm3 flask at 25°c and 98kPa pressure.


1
Expert's answer
2021-09-06T02:38:48-0400

According to ideal gas law:


"\frac{m}{M}×RT=PV"


Find the mass:

"m = \frac{(M×P×V)}{(R×T)}"


"V = 0.005 m^3"

"T = 298 K"

"P = 98000 Pa"

"R = 8.314 Jmol^{-1}K^{-1}"

"M = 71 g\/mol"


"m = \frac{(71×98000×0.005)}{(298×8.314) }= 14.04 g"